# HiLAB: A Hybrid Inverse-Design Framework

### v1

@@ Abstract
HiLAB (Hybrid inverse-design with Latent-space learning, Adjoint-based partial optimizations, and Bayesian optimization), the new paradigm for inverse design of nanophotonic structures presented in this paper, is a framework that combines early-terminated (partial) topological optimization (TO), a Vision Transformer–based variational autoencoder (VAE), and a Bayesian search strategy to address multi-functional photonic device design. By halting adjoint-driven TO well before full convergence—and randomizing physical parameters such as thickness and period—we generate a diverse set of freeform photonic configurations at a significantly reduced computational cost. These reasonably good structures are then embedded in a compact latent space by the VAE, enabling the Bayesian optimization to jointly optimize both the structural pattern (geometry) and physical hyperparameters (thickness, period) of metaphotonic devices. The learned latent space further facilitates (by transfer learning) for new objectives or constraints—such as fabrication tolerances—without necessitating a fresh round of partial TOs. Instead, the trained VAE can be reused, with modifications made only to the acquisition function in the Bayesian optimization to accommodate alternate design goals. Compared to conventional TO pipelines that typically converge to local optima, HiLAB systematically leverages uncertainty-guided sampling to explore clos-to-global optima or substantially improved local solutions while reducing electromagnetic simulation requirements. Even accounting for training overhead, HiLAB reduces the total number of full electromagnetic simulations by an order of magnitude, substantially accelerating the exploration of high-performance, fabrication-friendly optical devices. As a proof of concept, we apply HiLAB to the design of an achromatic beam deflector operating at red (660 nm), green (550 nm), and blue (470 nm) wavelengths. The resulting structure achieves balanced diffraction efficiencies—25.2% for red, 24.7% for green, and 24% for blue—while significantly mitigating chromatic aberrations, considerably surpassing all existing demonstrations. Beyond this specific application, HiLAB provides a powerful and generalizable strategy for enforcing multi-parameter constraints, achieving robust designs, and enabling rapid adaptation in designing next-generation nanophotonic and metamaterial devices.
@@
Inverse Design, Topology Optimization, Deep Learning, Bayesian Optimization, Achromatic Beam Deflector, Multi-wavelength Optics
@@ 1 Introduction
The development of new inverse-design techniques, empowered by topological optimization (TO), recently has revolutionized the way we engineer nanophotonic and metamaterial structures by algorithmically shaping device geometries to achieve specific optical responses  [1, 2, 3, 4, 5]. This paradigm has yielded significant breakthroughs in wavefront shaping [6, 7, 8], lensing and imaging [9, 10], nonlinear optics [11], and optical neural networks [12, 13, 14]. Among the most established inverse-design methods is gradient-based TO, which iteratively refines a continuous refractive-index distribution to maximize a desired figure of merit (FoM)[15, 16]. However, such methods are prone to convergence toward local optima, requiring hundreds of random restarts from different initial conditions to adequately sample the design landscape [17]. In addition, they typically fix crucial physical parameters—such as thickness or period—based on prior knowledge and/or assumptions about which physical modes or resonances might be supported. While this approach can be effective when parameters are either well understood or experimentally guided, it risks overlooking device geometries whose optimal performance only emerges under alternative or non-intuitive regimes. These drawbacks become especially pronounced in designing multi-functional (e.g., multi-wavelength) devices, demanding simultaneous optimization of performance across multiple spectral or polarization states. A second class of inverse-design approaches, shape optimization, more explicitly addresses physical parameters by describing devices in terms of relatively simple geometric primitives (e.g., elliptical or rectangular pillars) whose dimensions and positions can be systematically varied. Such approaches have successfully been applied to metasurface deflectors [7], polarization controllers [8], and mode converters [18], among others [19]. In addition, the simpler geometries produced by shape optimization are generally easier to fabricate than freeform layouts obtained from TO approaches. Nonetheless, restricting geometries to basic shapes can overlook complex freeform solutions vital for broadband or multi-functional tasks. Thus, while shape optimization systematically accounts for certain material or dimensional parameters, its limited geometric degrees of freedom often yield suboptimal designs when intricate 2D patterns are essential. A third category of inverse-design methods employs deep learning, leveraging neural networks trained on large datasets of device layouts and their corresponding optical responses [20, 21, 22]. This data-driven framework has inspired innovative demonstrations, such as inverse-designed color filters [23] and spectrally selective metasurfaces [20, 24]. In principle, once trained, these models can rapidly propose candidate structures without re-running exhaustive simulations. In practice, however, gathering thousands of labeled examples can be computationally expensive, especially since many randomly sampled structures may exhibit poor performance. Furthermore, most deep-learning pipelines rely on relatively simple device parameterizations—often shape-based—because fully freeform representations would require an excessively large number of design parameters, thus demanding even more extensive training data. From a physics perspective, such simplifications can forgo the opportunity to exploit mutual coupling or hybrid modes within complex freeform geometries, which often lead to higher performance. As a result, these methods may overlook richer design spaces that enable coupling intricate geometries with varying physical parameters, a limitation that grows more severe in multi-functional contexts.
@@
Compounding these conceptual hurdles are practical constraints related to electromagnetic (EM) solver speeds and robust device design. For instance, commercial finite-difference time-domain (FDTD) solvers (e.g., Ansys Lumerical) become prohibitively expensive for large-area devices with subwavelength resolution (e.g., millimeter-scale devices at visible wavelengths). Meanwhile, all listed inverse-design techniques, especially purely artificial intelligence (AI)-driven approaches, require large, high-quality training sets; yet random large-scale structures frequently fail to deliver meaningful performance, complicating dataset generation [25]. Finally, ensuring reliable device operation under inevitable fabrication imperfections remains particularly challenging for broadband or multi-functional designs.
@@
In response to these obstacles, we present in this paper HiLAB (Hybrid inverse-design with Latent-space learning, Adjoint-based partial optimizations, and Bayesian optimization), a framework that merges the versatility of TO with the search efficiency and dimensionality reduction offered by modern AI tools. Rather than performing fully converged TO at a fixed thickness or period of a metasurface, we truncate each TO run early and randomly vary key physical parameters, producing a compact dataset of reasonably good (though not optimal) freeform designs. We then train a variational autoencoder (VAE) that leverages a pre-trained Vision Transformer (ViT) for robust feature extraction, enabling it to learn a low-dimensional representation of these partial solutions—including fine-scale freeform details—while drastically reducing the computational cost of subsequent explorations. Finally, a Bayesian optimizer—requiring on the order of 300 evaluations—operates within this learned latent space, enabling multi-functional inverse design under practical constraints. As a concrete demonstration, our framework targets a multi-wavelength achromatic beam deflector that preserves identical steering angles for red, green, and blue visible light in a single-layer metasurface. By systematically sampling both freeform geometries and physical parameters, we uncover device configurations and physical modes that remain elusive in conventional pipelines. Inverse design of such an important device with conventional TO typically requires 200 iterations per run and 70 restarts to probe different basins of attraction, resulting in a total simulation budget of 14,000 forward EM solves. If fabrication robustness is considered, conventional robust TO methods typically require additional gradient evaluations for multiple variations of the nominal structure (such as eroded and dilated patterns) [29, 28, 6], thereby considerably increasing the standard computational workload. By contrast, our method requires only 1050 simulations to train a VAE model and 350 additional evaluations for Bayesian optimization (BO), achieving considerably better performance with just 1400 simulations—more than an order-of-magnitude reduction in cost [30], agnostic to the choice of underlying solver. Furthermore, the trained VAE can be easily repurposed for new objectives—such as fabrication tolerances—by simply adjusting the BO acquisition function. Consequently, the need to rerun partial TO is minimized, dramatically reducing the total number of EM simulations. Beyond this specific example, our hybrid approach—combining partial TO, VAE-based dimensionality reduction, and BO—shows significant promise for optimally designing the next generation of robust, large-scale multi-functional nanophotonic devices.
@@ 2 Methodology
2.1 Metphotonic Structure for Optimization
@@
where $n_{t}$ ($n_{i}$) represents the refractive index of transmitted (incident) medium; $\theta_{t}$ ($\theta_{i}$) denotes the transmitted (incident) angle, $\lambda$ is the operating wavelength, $\Lambda$ is the unit-cell period, and $m$ denotes the diffraction order. Fixing the period of the unit-cell at 5 µm along the x-axis under normal incidence yields a consistent deflection (or diffraction) angle of approximately 41.3° for all three wavelengths, effectively minimizing chromatic dispersion and enabling achromatic beam deflection. The period along the $y$-axis, $\Lambda_{y}$, is treated as a variable design parameter and jointly optimized alongside the device thicknesses (t1 and t2) and in-plane geometry as shown in Figure 1(a).
@@ 2.2 Overall Framework
Figure 1: Overview of the demonstrated HiLAB pipeline for multi-wavelength metasurface optimization. The metasurface consists of a patterned bilayer comprising SiO2 and TiO2, fabricated on top of a fused silica substrate. (a) Multiple partial TO runs under varied physical parameters, including $t_{1}$ (TiO2 thickness), $t_{2}$ (SiO2 thickness), and the unit-cell period $\Lambda_{y}$. The period along the $x$-direction, $\Lambda_{x}$, is fixed at 5 $\mu$m to achieve a 41.3 deflection angle. (b) Morphological data augmentation expands the library without additional EM simulations. (c) VAE-based dimensionality reduction (using a pre-trained ViT) learns a compact design manifold. (d) BO is performed jointly over the 8-dimensional geometry latent space and physical design parameters—including ($t_{1}$, $t_{2}$, and $\Lambda_{y}$) to identify metasurface layouts with optimal broadband performance. Each proposed parameter set is decoded into a structural design using the VAE decoder and evaluated through full-wave FDTD simulations. The resulting deflection efficiencies at the three design wavelengths (470 nm, 550 nm, and 660 nm) are used to compute a scalar FoM, defined as the worst-case efficiency across the spectrum: $\text{FoM}=\min\{\eta_{470},\eta_{550},\eta_{660}\}$. This conservative formulation ensures robust spectral performance and serves as the objective function value returned to the Bayesian optimizer. The pipeline incorporates post-decoding smoothing and binarization to ensure that final layouts comply with fabrication constraints such as minimum feature size and pattern co-registration. A 2D Principal Component Analysis (PCA) projection of the 11-dimensional sampled design vectors $x$ is shown to visualize optimizer performance and convergence.
In the third step (Figure 1(c)), we train a VAE—whose encoder uses a pretrained ViT—to learn a compact latent representation of the partially optimized structures. The ViT’s global attention mechanism readily accommodates two-dimensional (2D) metasurface patterns in an image-like form, while only the last few layers remain unfrozen to adapt to the metasurface dataset. This VAE projects each freeform topology into a low-dimensional latent space, preserving salient physics-driven features while omitting computationally expensive fine-scale degrees of freedom. By design, the decoder can regenerate high-resolution device layouts from latent vectors, effectively serving as a learned “reparameterization” that includes freeform geometry alongside any randomizable physical parameters (e.g., thickness and period). As a final step (Figure 1(d)), each layout undergoes Gaussian smoothing using a zero-mean Gaussian filter with a fixed standard deviation $\sigma=2.0$, followed by binarization with a threshold $\tau=0.6$. The Gaussian smoothing is applied via convolution with a 2D isotropic kernel, effectively suppressing subwavelength variations and small-scale noise that are difficult to fabricate. The thresholding operation then converts the smoothed layout into a binary pattern, ensuring compatibility with fabrication constraints and minimum feature size requirements. Once trained, the VAE’s decoder is embedded in a BO loop, wherein each candidate design is specified by a latent vector plus physical-parameter values as shown in Figure 1(d).
@@ 2.3 Partial TO and Data Augmentation
Our partial TO approach, involves truncating each adjoint-based optimization run, well before full convergence, at 35 iterations instead of the usual 150–450 [6, 28, 29, 32]. As shown in Figure 2(a),the principal design domain (The TiO2 layer) serves as the primary random design region and is discretized into a $256\times 128$ pixels along the $x$- and $y$-directions, respectively, where the refractive index is modulated during adjoint-based TO. As discussed in Section 2.1 and illustrated in Figure 1(a), our optimization goal is to maximize transmission efficiencies for selected diffraction orders at target wavelengths. At the end of optimization, the refractive indices of both TiO2 and SiO2 layers are binarized to their respective constituent materials resulting in a co-patterned, fabrication-ready bilayer structure. This ensures compatibility with standard lithographic processes and minimizes overlay errors during fabrication.
To promote smooth, manufacturable patterns and to guide convergence, we apply two regularization techniques at every iteration of the TO process: a cone filter and a progressive binarization function. The cone filter acts as a spatial smoothing operator, convolving the refractive index distribution with a circular kernel to suppress sub-wavelength noise and discontinuities. Its effective radius is updated at each iteration $i$ according to a logistic schedule:
@@
Figure 2: Evolution of a metasurface design across partial TO iterations 1, 18, 27, and 35. (a) The refractive-index profile progresses from near-random initialization (iteration 1) to a predominantly binarized pattern (iteration 35). (b) The corresponding deflection efficiencies at red (660 nm), green (550 nm), and blue (470 nm) wavelengths are plotted as functions of outgoing angle. The target (design) deflection angle is $41.3^{\circ}$ for all wavelengths, and each efficiency peak shifts progressively closer to this common angle over successive iterations. (c) Electric-field distributions in the $x\!z$-plane for the partially optimized structure (iteration 35) are shown for the three wavelengths. The black arrow highlights the deflected wavefront moving toward the desired output angle.
Finally, as shown in Figure 1(b), we augment each device layout by applying a suite of morphological transformations—specifically Gaussian smoothing, erosion, and dilation—to mimic the geometric transitions typically observed during the intermediate to final stages of gradient-based TO [31, 33]. This is motivated by the observation that in these later TO stages, the structures tend to evolve from noisy or fragmented configurations toward smoother, more regular patterns, as the algorithm refines features to maximize performance while improving fabricability. To emulate this progression, we apply Gaussian filters with varying standard deviations ($\sigma\in\{1.5,2.0,2.5\}$) where $\sigma$ is specified in pixel units, and controls the spatial extent of smoothing in the image domain, followed by morphological operations that simulate erosion (material removal) or dilation (feature expansion). These operations are systematically combined to produce variants such as Gaussian + erosion and Gaussian + dilation, in addition to standalone transformations, thereby capturing a broad range of refined topologies characteristic of late-stage optimization. This augmentation pipeline yields a diverse set of plausible designs without requiring any additional partial TO runs that require more extensive EM simulations. Staring with 30 partially optimized designs through TO, we generate different structured combinations of augmentation parameters and then fill out the dataset to 4000 samples using randomly selected transformation types and parameters. The resulting collection reflects realistic trajectories through the design space, encompassing both partially optimized devices and morphology-informed variants of likely high-performance solutions to train our VAE (Section 2.4). By embedding the morphological patterns characteristic of late-stage optimization, this augmentation strategy enhances model generalization and facilitates exploration of high-quality solutions in the downstream inverse design.

### v2

@@ Abstract
HiLAB (Hybrid inverse-design with Latent-space learning, Adjoint-based partial optimizations, and Bayesian optimization), the new paradigm for inverse design of nanophotonic structures presented in this paper, is a framework that combines early-terminated (partial) topological optimization (TO), a Vision Transformer–based variational autoencoder (VAE), and a Bayesian search strategy to address multifunctional photonic device design. By halting adjoint-driven TO well before full convergence—and randomizing physical parameters such as thickness and period—we generate a diverse set of freeform photonic configurations at a significantly reduced computational cost. These reasonably good structures are then embedded in a compact latent space by the VAE, enabling the Bayesian optimization to jointly optimize both the structural pattern (geometry) and physical hyperparameters (thickness, period) of metaphotonic devices. The learned latent space further facilitates (by transfer learning) for new objectives or constraints—such as fabrication tolerances—without necessitating a fresh round of partial TOs. Instead, the trained VAE can be reused, with modifications made only to the acquisition function in the Bayesian optimization to accommodate alternate design goals. Compared to conventional TO pipelines that typically converge to local optima, HiLAB systematically leverages uncertainty-guided sampling to explore close-to-global optima or substantially improved local solutions while reducing electromagnetic simulation requirements. Even accounting for training overhead, HiLAB reduces the total number of full electromagnetic simulations by an order of magnitude, substantially accelerating the exploration of high-performance, fabrication-friendly optical devices. As a proof of concept, we apply HiLAB to the design of an achromatic beam deflector operating at red (660 nm), green (550 nm), and blue (470 nm) wavelengths. The resulting structure achieves balanced diffraction efficiencies—25.2% for red, 24.7% for green, and 24% for blue—while significantly mitigating chromatic aberrations, considerably surpassing all existing demonstrations. Beyond this specific application, HiLAB provides a powerful and generalizable strategy for enforcing multi-parameter constraints, achieving robust designs, and enabling rapid adaptation in designing next-generation nanophotonic and metamaterial devices.
@@
Inverse Design, Deep Learning, Topological Optimization, Bayesian Optimization, Achromatic Beam Deflector, Multi-wavelength Optics
@@ 1 Introduction
The development of new inverse-design techniques, empowered by topological optimization (TO), has recently revolutionized the way we engineer nanophotonic and metamaterial structures by algorithmically shaping device geometries to achieve specific optical responses  [1, 2, 3, 4, 5]. This paradigm has yielded significant breakthroughs in wavefront shaping [6, 7, 8], lensing and imaging [9, 10], nonlinear optics [11], and optical neural networks [12, 13, 14]. Among the most established inverse-design methods is gradient-based TO, which iteratively refines a continuous refractive-index distribution to maximize a desired figure of merit (FoM)[5, 15]. However, such methods are prone to convergence toward local optima, requiring hundreds of random restarts from different initial conditions to adequately sample the design landscape [16]. In addition, they typically fix crucial physical parameters—such as thickness or period—based on prior knowledge and/or assumptions about which physical modes or resonances might be supported. While this approach can be effective when parameters are either well understood or experimentally guided, it risks overlooking device geometries whose optimal performance only emerges under alternative or non-intuitive regimes. These drawbacks become especially pronounced in designing multifunctional devices (e.g., multi-wavelength or multi-polarization) [17], demanding simultaneous optimization of performance across multiple spectral or polarization states. A second class of inverse-design approaches, shape optimization, more explicitly addresses physical parameters by describing devices in terms of relatively simple geometric primitives (e.g., elliptical or rectangular pillars) whose dimensions and positions can be systematically varied. Such approaches have successfully been applied to metasurface deflectors [7], polarization controllers [8], and mode converters [18], among others [19]. In addition, the simpler geometries produced by shape optimization are generally easier to fabricate than freeform layouts obtained from TO approaches. Nonetheless, restricting geometries to basic shapes can overlook complex freeform solutions vital for broadband or multifunctional tasks. Thus, while shape optimization systematically accounts for certain material or dimensional parameters, its limited geometric degrees of freedom often yield suboptimal designs when intricate 2D patterns are essential. A third category of inverse-design methods employs deep learning, leveraging neural networks trained on large datasets of device layouts and their corresponding optical responses [20, 21, 22]. This data-driven framework has inspired innovative demonstrations, such as inverse-designed color filters [23] and spectrally selective metasurfaces [20, 24]. In principle, once trained, these models can rapidly propose candidate structures without re-running exhaustive simulations. In practice, however, gathering thousands of labeled examples can be computationally expensive, especially since many randomly sampled structures may exhibit poor performance. Furthermore, most deep-learning pipelines rely on relatively simple device parameterizations—often shape-based—because fully freeform representations would require an excessively large number of design parameters, thus demanding even more extensive training data. From a physics perspective, such simplifications can forgo the opportunity to exploit mutual coupling or hybrid modes within complex freeform geometries, which often lead to higher performance. As a result, these methods may overlook richer design spaces that enable coupling intricate geometries with varying physical parameters, a limitation that grows more severe in multifunctional contexts.
@@
Compounding these conceptual hurdles are practical constraints related to electromagnetic (EM) solver speeds and robust device design. For instance, commercial finite-difference time-domain (FDTD) solvers (e.g., Ansys Lumerical) become prohibitively expensive for large-area devices with subwavelength resolution (e.g., millimeter-scale devices at visible wavelengths). Meanwhile, all listed inverse-design techniques, especially purely artificial intelligence (AI)-driven approaches, require large, high-quality training sets; yet random large-scale structures frequently fail to deliver meaningful performance, complicating dataset generation [25]. Finally, ensuring reliable device operation under inevitable fabrication imperfections remains particularly challenging for broadband or multifunctional designs.
@@
In response to these obstacles, we present in this paper HiLAB (Hybrid inverse-design with Latent-space learning, Adjoint-based partial optimizations, and Bayesian optimization), a framework that merges the versatility of TO with the search efficiency and dimensionality reduction offered by modern AI tools. Rather than performing fully converged TO at a fixed thickness or period of a metasurface, we truncate each TO run early and randomly vary key physical parameters, producing a compact dataset of reasonably good (though not optimal) freeform designs. We then train a variational autoencoder (VAE) that leverages a pre-trained Vision Transformer (ViT) for robust feature extraction, enabling it to learn a low-dimensional representation of these partial solutions—including fine-scale freeform details—while drastically reducing the computational cost of subsequent explorations. Finally, a Bayesian optimizer—requiring on the order of 300 evaluations—operates within this learned latent space, enabling multifunctional inverse design under practical constraints. As a concrete demonstration, our framework targets a multi-wavelength achromatic beam deflector that preserves identical steering angles for red, green, and blue visible light in a single-layer metasurface. By systematically sampling both freeform geometries and physical parameters, we uncover device configurations and physical modes that remain elusive in conventional pipelines. Inverse design of such an important device with conventional TO typically requires 200 iterations per run and 70 restarts to probe different basins of attraction, resulting in a total simulation budget of 14,000 forward EM solves. If fabrication robustness is considered, conventional robust TO methods typically require additional gradient evaluations for multiple variations of the nominal structure (such as eroded and dilated patterns) [29, 28, 6], thereby considerably increasing the standard computational workload. By contrast, our method requires only 1050 simulations to train a VAE model and 350 additional evaluations for Bayesian optimization (BO), achieving considerably better performance with just 1400 simulations—more than an order-of-magnitude reduction in cost [30], agnostic to the choice of underlying solver. Furthermore, the trained VAE can be easily repurposed for new objectives—such as fabrication tolerances—by simply adjusting the BO acquisition function. Consequently, the need to rerun partial TO is minimized, dramatically reducing the total number of EM simulations. Beyond this specific example, our hybrid approach—combining partial TO, VAE-based dimensionality reduction, and BO—shows significant promise for optimally designing the next generation of robust, large-scale multifunctional nanophotonic devices.
@@ 2 Methodology
2.1 Metaphotonic Structure for Optimization
@@
where $n_{t}$ ($n_{i}$) represents the refractive index of transmitted (incident) medium; $\theta_{t}$ ($\theta_{i}$) denotes the transmitted (incident) angle, $\lambda$ is the operating wavelength, $\Lambda$ is the unit-cell period, and $m$ denotes the diffraction order. Fixing the period of the unit-cell at 5 µm along the x-axis under normal incidence yields a consistent deflection (or diffraction) angle of approximately 41.3 for all three wavelengths, effectively minimizing chromatic dispersion and enabling achromatic beam deflection. The period along the $y$-axis, $\Lambda_{y}$, is treated as a variable design parameter and jointly optimized alongside the device thicknesses (t1 and t2) and in-plane geometry as shown in Figure 1(a).
@@ 2.2 Overall Framework
[…]
In the third step (Figure 1(c)), we train a VAE—whose encoder uses a pretrained ViT—to learn a compact latent representation of the partially optimized structures. The ViT’s global attention mechanism readily accommodates two-dimensional (2D) metasurface patterns in an image-like form, while only the last few layers remain unfrozen to adapt to the metasurface dataset. This VAE projects each freeform topology into a low-dimensional latent space, preserving salient physics-driven features while omitting computationally expensive fine-scale degrees of freedom. By design, the decoder can regenerate high-resolution device layouts from latent vectors, effectively serving as a learned “reparameterization” that includes freeform geometry alongside any randomizable physical parameters (e.g., thickness and period). As a final step (Figure 1(d)), each layout undergoes Gaussian smoothing using a zero-mean Gaussian filter with a fixed standard deviation $\sigma=2.0$, followed by binarization with a threshold $\tau=0.6$. The Gaussian smoothing is applied via convolution with a 2D isotropic kernel, effectively suppressing subwavelength variations and small-scale noise that are difficult to fabricate. The thresholding operation then converts the smoothed layout into a binary pattern, ensuring compatibility with fabrication constraints and minimum feature size requirements. Once trained, the VAE decoder is embedded in a BO loop, wherein each candidate design is specified by a latent vector plus physical-parameter values as shown in Figure 1(d).
@@ 2.3 Partial TO and Data Augmentation
Our partial TO approach, involves truncating each adjoint-based optimization run, well before full convergence, at 35 iterations instead of the usual 150–450 [6, 28, 29, 32]. The number of iterations is a method hyperparameter that balances design quality against simulation time during training-set generation. Our analysis (Supplementary Figure S1) indicates that at  30–40 iterations, both the design performance and the core topology stabilize. Continuing optimization beyond this point yields diminishing returns at substantial computational cost. As shown in Figure 2(a),the principal design domain (The TiO2 layer) serves as the primary random design region and is discretized into a $256\times 128$ pixels along the $x$- and $y$-directions, respectively, where the refractive index is modulated during adjoint-based TO. As discussed in Section 2.1 and illustrated in Figure 1(a), our optimization goal is to maximize transmission efficiencies for selected diffraction orders at target wavelengths. At the end of optimization, the refractive indices of both TiO2 and SiO2 layers are binarized to their respective constituent materials resulting in a co-patterned, fabrication-ready bilayer structure. This ensures compatibility with standard lithographic processes and minimizes overlay errors during fabrication.
To promote smooth, manufacturable patterns and to guide convergence, we apply two regularization techniques at every iteration of the TO process: a cone filter and a progressive binarization function [6, 16, 33]. The cone filter acts as a spatial smoothing operator, convolving the refractive index distribution with a circular kernel to suppress sub-wavelength noise and discontinuities. Its effective radius is updated at each iteration $i$ according to a logistic schedule:
@@
Figure 2: Evolution of a metasurface design across partial TO iterations 1, 18, 27, and 35. (a) The refractive-index profile progresses from near-random initialization (iteration 1) to a predominantly binarized pattern (iteration 35). (b) The corresponding deflection efficiencies at red (660 nm), green (550 nm), and blue (470 nm) wavelengths are plotted as functions of outgoing angle. The target (design) deflection angle is $41.3^{\circ}$ for all wavelengths, and each efficiency peak shifts progressively closer to this common angle over successive iterations. (c) Electric-field distributions in the xzx\!z-plane for the partially optimized structure (iteration 35) are shown for the three wavelengths. The black arrow highlights the deflected wavefront moving toward the desired output angle.
Finally, as shown in Figure 1(b), we augment each device layout by applying a suite of morphological transformations—specifically Gaussian smoothing, erosion, and dilation—to mimic the geometric transitions typically observed during the intermediate to final stages of gradient-based TO [31, 33]. This is motivated by the observation that in these later TO stages, the structures tend to evolve from noisy or fragmented configurations toward smoother, more regular patterns, as the algorithm refines features to maximize performance while improving fabricability. To emulate this progression, we apply Gaussian filters with varying standard deviations ($\sigma\in\{1.5,2.0,2.5\}$) where $\sigma$ is specified in pixel units, and controls the spatial extent of smoothing in the image domain, followed by morphological operations that simulate erosion (material removal) or dilation (feature expansion). These operations are systematically combined to produce variants such as Gaussian + erosion and Gaussian + dilation, in addition to standalone transformations, thereby capturing a broad range of refined topologies characteristic of late-stage optimization. This augmentation pipeline yields a diverse set of plausible designs without requiring any additional partial TO runs that require more extensive EM simulations. Starting with 30 partially optimized designs through TO, we generate different structured combinations of augmentation parameters and then fill out the dataset to 4000 samples using randomly selected transformation types and parameters. The resulting collection reflects realistic trajectories through the design space, encompassing both partially optimized devices and morphology-informed variants of likely high-performance solutions to train our VAE (Section 2.4). By embedding the morphological patterns characteristic of late-stage optimization, this augmentation strategy enhances model generalization and facilitates exploration of high-quality solutions in the downstream inverse design.